OpenStudy (anonymous):

Where is the vertex of the parabola? y = x2 + 2x + 3 is it above the x-axis below the x-axis on the x-axis on the y-axis

OpenStudy (anonymous):

You need to find the x coordinate of the vertex by using x = -b/2a so x = -2/2 x = -1 so when x = -1 you find the vertex. Put in x = -1 and that will give you the y coordinate